OpenStudy (redneckoutlaw):

What is the exact value of the square root of the quantity 99 over 55.? Simplify if possible. Choose one answer. a. 3 the square root of 5 all over 5. b. 3 the square root of 11 all over 5. c. 11 the square root of 3 all over 5. d. 5 the square root of 3 all over 5.

OpenStudy (redneckoutlaw):

@triciaal

OpenStudy (anonymous):

do you know how to simplify square roots??

OpenStudy (cwrw238):

list the prime factors of 99 and 55 99 = 3*3*11 55 = 5*11 can you see a perfect square amongst the above?

OpenStudy (cwrw238):

also if we divide 99 by 55 can you see a common factor/